OpenStudy (anonymous):

Solve the system by substitution. 2x-y+z=-4 z=5 -2x+3y-z=-10 a.) (-8, 7, 5) b.) (–8, –7, 5) c.) (8, –7, 5) d.)(–8, –7, –5)

OpenStudy (anonymous):

@amorfide Please help.. I can't do this

OpenStudy (anonymous):

Still there?

OpenStudy (anonymous):

Ok. This is actually simple. We already know Z = 5. So substitute that into any of the two remaining equations. Ill go with 2x - y + 2 = -4.

OpenStudy (anonymous):

I mean. 2x - y + z = -4 Sorry about that.

OpenStudy (anonymous):

So now we have 2x -

OpenStudy (anonymous):

2x - y + 5 = -4 --> -y = -2x -5 - 4 or otherwise y = 2x +5 + 4

OpenStudy (anonymous):

Now substitute that into -2x + 3y - z = -10. So -2x + 3(2x + 5 + 4) - 5 = -10 Now. Itll end up being -2x + 6x + 15 + 12 - 5 = -10. Which will then become 4x + 15 + 12 - 5 = -10. Hopefully you know what to do after that. x = -8.

OpenStudy (anonymous):

So. We have x = -8, and z = 5. Substitute into any equation. Ill use 2x - y + z = -4. Which will be -16 - y + 5 = -4. --> y = -16 + 5 + 4. SO Y = -7. X = -8 and Z = 5. Hope you get it now. :)

OpenStudy (anonymous):

-waiting patiently for my best answer medal- lol